OpenStudy (anonymous):

The lengths of the sides in a right triangle form three consecutive terms of a geometric sequence. Find the common ratio of the sequence. There are two distinct answers.

OpenStudy (anonymous):

A. 10.0 cm B. 34.6 cm C. 85.1 cm D. 120.0 cm which on would it be ?>

OpenStudy (anonymous):

i need to find the common ratio not a length

OpenStudy (anonymous):

Ok. so let the required ratio be R. Then is it true that - the sides will be a, aR, aR^2 with aR^2 as the largest side i.e. hypotenuse?

OpenStudy (anonymous):

i had a/r,a,ar but sure

OpenStudy (anonymous):

@luckythebest

OpenStudy (anonymous):

then is it true that - a^2 + a^2R^2 = a^2R^4 By pythagorean theorem?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

soooo

OpenStudy (anonymous):

is it further true that - a^2(1+R^2) = a^2R4 => 1+R^2 = R^4 ?

OpenStudy (anonymous):

Let the side lengths be \(1, a, a^2\) for some nonnegative number \(a\). If \(a > 1\), then \(a^2\) is the longest side so \((a^2)^2=(1)^2+(a)^2\). If \(0 < a < 1\), then \(1\) is the longest side so \((1)^2=(a)^2+(a^2)^2\). We can omit the \(a=1\) case since that would be an equilateral triangle. This should give you enough information to come up with the two distinct answers.

OpenStudy (anonymous):

@yakeyglee who said 0<a<1

OpenStudy (anonymous):

and yes lucky

OpenStudy (anonymous):

Next - Let R^2 = K Then is it true that - 1+K = K^2 ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

We have - K^2 - K - 1 = 0 Compare with ax^2 + bx + c = 0 a = 1, b = -1, c = -1 Find b^2-4ac and substitute it in the quadratic formula. Solve for K, and tell me the answer.

OpenStudy (anonymous):

If we assume \(0<a<1\) that is one of the two cases. It affects how you treat the computation.

OpenStudy (anonymous):

@yakeyglee if you substitute ^ in the quadratic formula you anyway get 2 distinct answers; example x+a and x-a, isn't it?

OpenStudy (anonymous):

(1(+-)sqrt5)/2

OpenStudy (anonymous):

Ok. so K = 1 +/- Sqrt 5/2 But we told earlier that K = R^2 Hence R^2 = 1 +/- sqrt 5/2 R = Sqrt of R^2. The answer is a bit complex but that is how its done. :(

OpenStudy (anonymous):

so r=sqrt((1+/- sqrt5)/2)

OpenStudy (anonymous):

:like:

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i mean its correcto :)

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

If you want further simplification sqrt 5 = 2.236 then R = (1+2.236/1.414) => sqrt of 3.236 / 1.414 (1.414 is sqrt of 2)

OpenStudy (anonymous):

@luckythebest I didn't use any x in my computation... i don't understand what you are talking about. And you will get multiple values for \(a\), but I bet only one of them satisfies the initial assumption that was made to derive the equation.

OpenStudy (anonymous):

@yakeyglee what i meant to say was that if you use the quadratic equation formula anyway you will get two distinct answers.